a cart rolls 2 m to the right then rolls back 1 m to the left . what is the displacement of the cart from the initial to the final position


1
Expert's answer
2021-04-19T17:06:39-0400

Assume that to the right is a positive displacement and to the left is negative.

The total displacement is the sum of displacements:


"d=d_1+d_2=2+(-1)=1\text{ m}."